(-1,1),(2,5)
how could you get an equation from graphing the coordinates?

Found 2 solutions by mathie123, ewatrrr:
Answer by mathie123(224)

Does the question say that these two points lie on a line (that would be what I would guess but just making sure then I can help you further)

Answer by ewatrrr(24785)

 
Hi,
how could you get an equation from graphing the coordinates?(-1,1),(2,5)
1. Plot Points and connect with line y.
Note: y = mx + b, where m is the slope and b the y-intercept
2)Determine where the line crosses the y-axis: 2 1/3
3) Determine the slope:
going from (-1,1) to (2,5) one goes 4up and 3 over..Slope = 4/3
y = (4/3)x + 7/3 or in Standard form: 4x-3y = -7